HIGH RESOLUTION FTIR SPECTRA OF METHYL NITRITE AND ASSIGNMENT OF THE ν_{8} FUNDA-MENTAL OF THE CIS CONFORMER

\underline{LISA M. GOSS}, Department of Chemistry, Idaho State University, Campus Box 8023, Pocatello ID, 83209; THOMAS A. BLAKE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99352.

The infrared spectrum of gas phase methyl nitrite (CH_{3}ONO) has been recorded at high resolution between 600 and 4000 cm^{−1}. The conformational doublets due to the cis and trans conformers are clearly separated in a number of the fundamentals. Previous work in the microwave has established that the barrier to rotation of the methyl group in the trans form is very low^{a} and the infrared bands observed in this study are very congested. The barrier in the cis conformer is much higher^{b}. The ν_{8} at 841 cm^{−1} was assigned by previous workers to the N-O stretch of the cis conformer^{c} and is partially overlapped by the ν_{8} of the trans conformer. In this study, spectra taken at 0.0015 cm^{−1} resolution on the PNNL IFS120/HR in a 20 cm cell were used to assign more than 800 A-type rovibrational lines of the cis ν_{8} between 815 and 865 cm^{−1}.

aTurner, P. H., et al., J. Phys. Chem. 83 (11), 1473 1979.

bGhosh, P. N., et al., Chem. Phys. 53 (1-2), 39 1980.

cStidham, H. D., et al. J. Raman Spec. 21 (1-2), 615 1990; da Silva, J. B. P., et al. J. Molec. Struct. 375 (1/2), 153 1996; and references therein.